A student has 6 different notebooks and decided to give 5 of them to a friend. How many ways does he have to give away these notebooks?

1
Expert's answer
2022-03-29T12:14:07-0400

As the order of notebooks does not matter, we'll count the number of combinations without repetition:

"N=\begin{pmatrix} 6\\ 5\end{pmatrix}=\\
=\cfrac{6! } {5! \cdot(6-5)! }=6."